OpenStudy (anonymous):

simplify each expression. write each answer in scientific notation. (2x10^3)(3x10^2)

OpenStudy (anonymous):

(2x10^3)= (2x1000)(3x100) = (2000x300)= 600000= 6X10^5

OpenStudy (anonymous):

6.0 X 10^5

OpenStudy (ohmaigosshh):

(2x10^3) = 2,000 Sci. Notation: 2x10^3and (3x10^2) = 300 Sci. Noatation: 3x10^2

OpenStudy (anonymous):

thank you very much!